Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 0. Let reserves = 200. Consider the basic two-period model. Assume that r increases from 10% to 20%. What happens to price and quantity in the present?

Price and quantity both increase

Price increases and quantity decreases

Price decreases and quantity increases

Price and quantity both decrease
Answer #1

Answer option 3)

Q = 250-P

At eqm,

250-Q1 = (250-Q2)/(1+r)

Q1+Q2= 200

Put Q2= 200-Q1

If r = .1

Then Q1*= 107.143

If r= .2

Then Q1"= 113.636

So present Quantity rises, so Price decrease